OpenStudy (anonymous):

se the Distributive Property and simplify the expression. 9 + 4(2x – 1) + 8 A. 2x + 20 B. 8x + 16 C. 8x + 13 D. 6x + 20

OpenStudy (anonymous):

8x+13

OpenStudy (anonymous):

i think maybe be c

OpenStudy (igreen):

9 + 4(2x - 1) + 8 You have to distribute the 4 into the parenthesis by multiplying it to every term in the parenthesis. 4 * 2x = 8x 4 * -1 = -4 So: 9 + 8x - 4 + 8 Add: 9 + 8x + 4 8x + 13

OpenStudy (rianiscool):

c 8x+13

OpenStudy (igreen):

@heavenfun13 @godnessisfone @rianiscool You're not supposed to give out direct answers without any explanation.

OpenStudy (rianiscool):

se what agree said 9 + 4(2x - 1) + 8 You have to distribute the 4 into the parenthesis by multiplying it to every term in the parenthesis. 4 * 2x = 8x 4 * -1 = -4 So: 9 + 8x - 4 + 8 Add: 9 + 8x + 4 8x + 13

OpenStudy (anonymous):

9 + 4(2x - 1) + 8 You have to distribute the 4 into the parenthesis by multiplying it to every term in the parenthesis. 4 * 2x = 8x 4 * -1 = -4 So: 9 + 8x - 4 + 8 Add: 9 + 8x + 4 8x + 13